OpenStudy (anonymous):

The product of (a − b)(a − b) is a^2 − b^2. Sometimes Always Never

OpenStudy (lgbasallote):

this is a rule...the difference of two squares \[a^2 - b^2 = (a+b)(a-b)\] you can prove it by using FOIL \[(a+b)(a-b) = a^2 - ab + ab + b^2 \implies a^2 -b^2\]

OpenStudy (anonymous):

so its always

OpenStudy (lgbasallote):

Yes.

OpenStudy (anonymous):

Thanks !

OpenStudy (lgbasallote):

welcome